Find the inverse of the given function. Upload solution in next number.

Let f:R→R , f(x)= 4x + 3

a. f-1 (y)= 

b. f-1 (35) = 

c. f-1 (-9) = 


1
Expert's answer
2022-04-12T09:57:02-0400
  • "Let f ^{\n\u22121}\n (x)=y \n\u21d2 x=f(y)\n\u21d2 x=4y+3\n\u21d2 x\u22123=4y\n\u21d2 y= \\frac{\n\nx\u22123\n\u200b}{4}"

b."f^{-1}(35)=\frac{35-3}{4}=8"

c. "f^{-1}(-9)=\frac{-9-3}{4}=-3"